Question 7


(7.1) (Find a point-normal form of the equation of the plane passing

through P = (1,2,−3) and having n =< 2,−1,2 > as a normal.


(7.2) Determine in each case whether the given planes are parallel or

perpendicular: (a) x +y +3z +10=0andx +2y −z =1,


(b)3x −2y +z −6=0 and 4x +2y −4z =0, (c)3x +y +z −1=0 and−x +2y

+z+3=0,


(d)x −3y +z+1=0 and 3x −4y +z−1=0.


1
Expert's answer
2021-07-16T07:44:57-0400


(7.1)

Point-normal form of the plane is



"2(x \u2212 1) -(y \u2212 2) + 2(z +3) = 0."

We can also write this as



"2x-y+2z=-6"



(7.2)

(a). x+y+3z+10=0 and x+2y-z=1



"\vec n_1=\langle1,1,3\rangle, \vec n_2=\langle1,2,-1\rangle""\dfrac{1}{1}\not=\dfrac{2}{1}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.




"\vec n_1\cdot\vec n _2=1(1)+1(2)+3(-1)=0"

The vectors "\vec n_1" and "\vec n_2" are orthogonal.

Hence, the given planes are perpendicular.


(b). 3x-2y+z-6=0 and 4x+2y-4z=0



"\vec n_1=\langle3,-2,1\rangle, \vec n_2=\langle4,2,-4\rangle""\dfrac{4}{3}\not=\dfrac{2}{-2}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.




"\vec n_1\cdot\vec n _2=3(4)+(-2)(2)+1(-4)=4\not=0"

The vectors "\vec n_1" and "\vec n_2" are not orthogonal.

Hence, the given planes are neither parallel nor perpendicular.


(c). 3x+y+z-1=0 and -x+2y+z+3=0



"\vec n_1=\langle3,1,1\rangle, \vec n_2=\langle-1,2,1\rangle""\dfrac{-1}{3}\not=\dfrac{2}{1}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.




"\vec n_1\cdot\vec n _2=3(-1)+1(2)+1(1)=0"

The vectors "\vec n_1" and "\vec n_2" are orthogonal.

Hence, the given planes are perpendicular.


(d). x-3y+z+1=0 and 3x-4y+z-1=0



"\vec n_1=\langle1,-3,1\rangle, \vec n_2=\langle3,-4,1\rangle""\dfrac{3}{1}\not=\dfrac{-4}{-3}"

The vectors "\vec n_1" and "\vec n_2" are not collinear.




"\vec n_1\cdot\vec n _2=1(3)+(-3)(-4)+1(1)=16\not=0"

The vectors "\vec n_1" and "\vec n_2" are not orthogonal.

Hence, the given planes are neither parallel nor perpendicular.